# On the Provable Suboptimality of Momentum SGD in Nonstationary Stochastic Optimization

### v2

@@ D.1 Conditional Orlicz norms
Theorem D.1 (Basic properties).
@@
Lemma D.1 (Conditional “sub-Gaussian $\times$ sub-Gaussian $\Rightarrow$ sub-exponential”).
@@
Lemma D.2 (Monotonicity and conditioning rules).
  1. (i)
@@
Lemma D.3 (A frequently used specialization).
Let $\mathcal{G}_{1}\subseteq\mathcal{G}_{2}$. If $\|X\mid\mathcal{G}_{1}\|_{\Psi_{2}}\leq\lambda$ a.s. and $\|\xi\mid\mathcal{G}_{2}\|_{\Psi_{2}}\leq K$ a.s. for constants $\lambda,K$, then
$\|\xi X\mid\mathcal{G}_{1}\|_{\Psi_{1}}~\leq~K\lambda\qquad\text{a.s.}$
@@
By Lemma˜D.2(i), $\|\xi\mid\mathcal{G}_{1}\|_{\Psi_{2}}\leq K$ a.s. Apply Lemma˜D.1 (in the scalar case) conditioned on $\mathcal{G}_{1}$ to obtain $\|\xi X\mid\mathcal{G}_{1}\|_{\Psi_{1}}\leq\|\xi\mid\mathcal{G}_{1}\|_{\Psi_{2}}\cdot\|X\mid\mathcal{G}_{1}\|_{\Psi_{2}}\leq K\lambda$ a.s. ∎
Lemma D.4 (Conditional $\Psi_{2}$ control implies conditional second moments).
@@
D.2 Martingale Concentration Inequalities
Lemma D.5 (Bernstein inequality for sub-exponential martingale differences).
@@ D.2 Martingale Concentration Inequalities
$\mathbb{P}\left(\sum_{t=1}^{T}Z_{t}\geq s\right)=\mathbb{P}\left(\exp\Bigl(\lambda\sum_{t=1}^{T}Z_{t}\Bigr)\geq e^{\lambda s}\right)\leq e^{-\lambda s}\,\mathbb{E}\exp\Bigl(\lambda\sum_{t=1}^{T}Z_{t}\Bigr).$ (D.1)
@@
$\mathbb{E}\exp\Bigl(\lambda\sum_{t=1}^{T}Z_{t}\Bigr)=\mathbb{E}e^{\lambda S_{T}}\leq\exp\left(C\lambda^{2}\sum_{t=1}^{T}K_{t}^{2}\right).$ (D.2)
Combining (D.1) and (D.2) gives, for all admissible $\lambda$,
@@
D.3 Martingale Theory
Lemma D.6 (Optional stopping / sampling theorem).
@@ D.3 Martingale Theory
D.4 Gaussian Concentration Inequalities
Lemma D.7 (Gaussian small-ball bound from a variance lower bound).
Let $X$ be a real-valued Gaussian random variable with $\mathrm{Var}(X)\geq v_{\min}>0$. Then for every $r>0$,
$\mathbb{P}(|X|\leq r)\;\leq\;\sqrt{\frac{2}{\pi}}\;\frac{r}{\sqrt{v_{\min}}}.$ (D.3)
Moreover, the bound is sharp in the sense that equality holds in (D.3) whenever $X\sim\mathcal{N}(0,v_{\min})$ and $r\downarrow 0$ (up to the first-order Taylor expansion).
Proof.
Since $X$ is Gaussian, there exist $m\in\mathbb{R}$ and $v>0$ such that $X\sim\mathcal{N}(m,v)$. By assumption, $v=\mathrm{Var}(X)\geq v_{\min}$. Let $f_{m,v}$ denote the Lebesgue density of $\mathcal{N}(m,v)$, i.e.,
$f_{m,v}(x):=\frac{1}{\sqrt{2\pi v}}\exp\!\left(-\frac{(x-m)^{2}}{2v}\right),\qquad x\in\mathbb{R}.$ (D.4)
For any Borel set $A\subseteq\mathbb{R}$, the probability of $X$ falling in $A$ is given by
$\mathbb{P}(X\in A)=\int_{A}f_{m,v}(x)\,dx,$ (D.5)
where the integral is with respect to Lebesgue measure. We apply (D.5) to the interval A=[r,r]A=[-r,r]:
$\mathbb{P}(|X|\leq r)=\mathbb{P}(X\in[-r,r])=\int_{-r}^{r}f_{m,v}(x)\,dx.$ (D.6)
The key observation is that (D.6) can be upper bounded by the length of the interval times the supremum of the density on that interval:
$\int_{-r}^{r}f_{m,v}(x)\,dx\;\leq\;\int_{-r}^{r}\sup_{y\in\mathbb{R}}f_{m,v}(y)\,dx=(2r)\,\sup_{y\in\mathbb{R}}f_{m,v}(y).$ (D.7)
Thus, it suffices to upper bound $\sup_{y\in\mathbb{R}}f_{m,v}(y)$ in terms of $v_{\min}$. From (D.4), the exponential factor satisfies
$0<\exp\!\left(-\frac{(x-m)^{2}}{2v}\right)\leq 1\qquad\text{for all }x\in\mathbb{R},$
with equality if and only if $x=m$. Therefore,
$\sup_{y\in\mathbb{R}}f_{m,v}(y)=f_{m,v}(m)=\frac{1}{\sqrt{2\pi v}}.$ (D.8)
Since $v\geq v_{\min}$, we have $\frac{1}{\sqrt{v}}\leq\frac{1}{\sqrt{v_{\min}}}$ and hence
$\sup_{y\in\mathbb{R}}f_{m,v}(y)=\frac{1}{\sqrt{2\pi v}}\;\leq\;\frac{1}{\sqrt{2\pi v_{\min}}}.$ (D.9)
Substituting (D.9) into (D.7) and then into (D.6) yields
$\mathbb{P}(|X|\leq r)\leq(2r)\cdot\frac{1}{\sqrt{2\pi v_{\min}}}=\sqrt{\frac{2}{\pi}}\;\frac{r}{\sqrt{v_{\min}}},$
which is exactly (D.3). For $X\sim\mathcal{N}(0,v_{\min})$, the density at 0 is $(2\pi v_{\min})^{-1/2}$, so as $r\downarrow 0$,
$\mathbb{P}(|X|\leq r)=\int_{-r}^{r}f_{0,v_{\min}}(x)\,dx=2rf_{0,v_{\min}}(0)+o(r)=\sqrt{\frac{2}{\pi}}\;\frac{r}{\sqrt{v_{\min}}}+o(r),$
showing that the constant $\sqrt{2/\pi}$ is tight to first order. ∎
D.5 Information Theory
Lemma D.8 (Mutual information bounded by average pairwise KL).
@@ D.5 Information Theory
$I(U;Z)\;=\;\frac{1}{M}\sum_{u\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}\,\|\,\bar{P})\;\leq\;\frac{1}{M^{2}}\sum_{u,v\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}\,\|\,P_{v}).$ (D.10)
Proof of Lemma˜D.8.
@@
$I(U;Z)=\sum_{u\in\mathcal{U}}\frac{1}{M}\int\log\!\left(\frac{dP_{u}}{d\bar{P}}(z)\right)\,dP_{u}(z)=\frac{1}{M}\sum_{u\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}\,\|\,\bar{P}),$ (D.11)
@@
$\log\!\Big(\frac{1}{M}\sum_{v\in\mathcal{U}}p_{v}(z)\Big)\;\geq\;\frac{1}{M}\sum_{v\in\mathcal{U}}\log p_{v}(z).$ (D.12)
@@
$=\frac{1}{M}\sum_{v\in\mathcal{U}}\int\big(\log p_{u}(z)-\log p_{v}(z)\big)\,p_{u}(z)\,dz\;=\;\frac{1}{M}\sum_{v\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}\,\|\,P_{v}).$ (D.13)
Averaging (D.13) over $u\sim\mathrm{Unif}(\mathcal{U})$ and using (D.11) yields

### v3

@@ G.1 Conditional Orlicz norms
Theorem G.1 (Basic properties).
@@
Lemma G.1 (Conditional “sub-Gaussian $\times$ sub-Gaussian $\Rightarrow$ sub-exponential”).
@@
Lemma G.2 (Monotonicity and conditioning).
We have the following rules for conditioning and monotonicity:
  1. 1.
@@
Lemma G.3.
Let $\mathcal{G}_{1}\subseteq\mathcal{G}_{2}$. If $\|X\mid\mathcal{G}_{1}\|_{\Psi_{2}}\leq\lambda$ a.s. and $\|\bm{\xi}\mid\mathcal{G}_{2}\|_{\Psi_{2}}\leq K$ a.s. for constants $\lambda,K$, then
$\|\bm{\xi}X\mid\mathcal{G}_{1}\|_{\Psi_{1}}~\leq~K\lambda\qquad\text{a.s.}$
@@
By Lemma˜G.2(i), $\|\bm{\xi}\mid\mathcal{G}_{1}\|_{\Psi_{2}}\leq K$ a.s. Apply Lemma˜G.1 (in the scalar case) conditioned on $\mathcal{G}_{1}$ to obtain $\|\bm{\xi}X\mid\mathcal{G}_{1}\|_{\Psi_{1}}\leq\|\bm{\xi}\mid\mathcal{G}_{1}\|_{\Psi_{2}}\cdot\|X\mid\mathcal{G}_{1}\|_{\Psi_{2}}\leq K\lambda$ a.s. ∎
Lemma G.4 (Conditional $\Psi_{2}$ control implies conditional second moments).
@@
G.2 Martingale Concentration Inequalities
Lemma G.5 (Bernstein inequality for sub-exponential martingale differences).
@@ G.2 Martingale Concentration Inequalities
$\mathbb{P}\left(\sum_{t=1}^{T}Z_{t}\geq s\right)=\mathbb{P}\left(\exp\Bigl(\lambda\sum_{t=1}^{T}Z_{t}\Bigr)\geq e^{\lambda s}\right)\leq e^{-\lambda s}\,\mathbb{E}\exp\Bigl(\lambda\sum_{t=1}^{T}Z_{t}\Bigr).$ (244)
@@
$\mathbb{E}\exp\Bigl(\lambda\sum_{t=1}^{T}Z_{t}\Bigr)=\mathbb{E}e^{\lambda S_{T}}\leq\exp\left(C\lambda^{2}\sum_{t=1}^{T}K_{t}^{2}\right).$ (245)
Combining (244) and (245) gives, for all admissible $\lambda$,
@@
G.3 Optional stopping
Lemma G.6 (Optional stopping / sampling theorem).
@@ G.3 Optional stopping
G.4 Information Theory
Lemma G.7 (Mutual information bounded by average pairwise KL).
@@ G.4 Information Theory
$I(U;Z)\;=\;\frac{1}{M}\sum_{u\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}\,\|\,\bar{P})\;\leq\;\frac{1}{M^{2}}\sum_{u,v\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}\,\|\,P_{v}).$ (246)
Proof of Lemma˜G.7.
@@
$I(U;Z)=\sum_{u\in\mathcal{U}}\frac{1}{M}\int\log\!\left(\frac{dP_{u}}{d\bar{P}}(z)\right)\,dP_{u}(z)=\frac{1}{M}\sum_{u\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}\,\|\,\bar{P}),$ (247)
@@
$\log\!\Big(\frac{1}{M}\sum_{v\in\mathcal{U}}p_{v}(z)\Big)\;\geq\;\frac{1}{M}\sum_{v\in\mathcal{U}}\log p_{v}(z).$ (248)
@@
$=\frac{1}{M}\sum_{v\in\mathcal{U}}\int\big(\log p_{u}(z)-\log p_{v}(z)\big)\,p_{u}(z)\,dz\;=\;\frac{1}{M}\sum_{v\in\mathcal{U}}D_{\mathrm{KL}}(P_{u}\,\|\,P_{v}).$ (249)
Averaging (249) over $u\sim\mathrm{Unif}(\mathcal{U})$ and using (247) yields